Waller, Inc., is trying to determine its cost of debt. The firm has a debt issue outstanding with 13 years to maturity that is quoted at 104 percent of face value. The issue makes semiannual payments and has an embedded cost of 11 percent annually.

(a) What is the company's pretax cost of debt? (Do not round your intermediate calculations.)

11.20%

10.43%

9.91%

10.95%

10.85%

(b) If the tax rate is 35 percent, what is the aftertax cost of debt? (Do not round your intermediate calculations.)

6.44%

7.05%

5.90%

6.78%

7.12%
Answer #1

1.
Using financial calculator
N=13*2=26
PMT=11%*1000/2=55
PV=-104%*1000=-1040
FV=1000
CPT I/Y=5.22%

Pre-tax cost of debt=5.22%*2=10.43%

2.
After-tax cost of debt=Pre-tax cost of debt*(1-tax rate)=10.43%*(1-35%)=6.78%